# Essential difference between 2D and 3D from the perspective of real-space renormalization group

### v2

@@ Abstract
Block tensor in 3D.—The advantage of TNRG in 2D makes it a promising candidate as a 3D real-space RG. One thing to note is that, without the viewpoint of entanglement entropy and its area laws, whether the block-tensor RG in 3D is better or worse than that in 2D is not straightforward. Recall that in Wilson’s momentum-shell RG, systems in dimensions $d>4$ are easy, where mean-field treatment of Landau-Ginzburg field is enough; but in $d<4$, perturbative $\epsilon$-expansion is necessary and $d=2$ is harder than $d=3$ since $\epsilon=4-d$ is larger. Interestingly, numerical experiments [25, 26] would seemingly imply that block-tensor RG in 3D improves [RG flow condition] compared to 2D. However, the truncation effects, as a cause of those observations, become more transparent through the lens of entanglement entropy.
@@
Figure 1: (Upper left) A tensor $A^{\text{EDL}}(e)$ with an EDL structure is the tensor product of 12 copies of edge-matrix $e$ located at 12 edges of a cube block; every 4 matrix-indices on the same face are grouped together to be a thicker leg of the 6-leg $A^{\text{EDL}}(e)$ (Lower right) A block-tensor RG with rescaling factor $b=2$. A total of $12\times 8=96$ copies of edge-matrix $e$ form 3 types: (i) Around each of the 6 axes inside the block, 4 copies of $e$ form a loop and becomes a number, so there are $4\times 6$ copies; (ii) On each of 6 faces, copies of $e$ form 4 entanglement pairs inside the face, so there are $2\times 4\times 6$ copies; (iii) The remaining $2\times 12$ copies sit on the 12 edges of the larger block, each edge having 2 copies side by side. Only type-(iii) copies survive the exact block-tensor RG; the other two types become an overall multiplication factor. The coarser tensor has an EDL structure built from $e\otimes e$: $A^{\prime}=A^{\text{EDL}}(e\otimes e)$.
@@
The assessment of the block-tensor RG in 3D parallels its analysis in 2D. We have a 6-leg tensor $A_{\vec{s}_{x+},\vec{s}_{x-},\vec{s}_{y+},\vec{s}_{y-},\vec{s}_{z+},\vec{s}_{z-}}$ (for example, $x+$ denotes the leg along the positive $x$ axis), corresponding to a cubic block of spins. The TNRG map in Eq. (4) satisfies $Z_{N}(A)=Z_{N/b^{3}}(A^{\prime})$. For a large block with linear size $L$, the area law of 3D quantum field theories says that the entanglement entropy of a subsystem on a face of the cube, $S(\rho_{\vec{s}_{x+}})$ for example, scales like [27, 28]
@@
where $\alpha$ contains microscopic details and $F$ encodes universal properties of the system. This entropy should not be confused with the mutual information between the inside and the outside of the cube mediated through the face. Rather, it is from the correlations between the two surface areas mediated through the edge region. The area law in Eq. (5) reveals a lot about 3D block-tensor transformation. The growth of $S(\rho_{\vec{s}_{x+}})$ with the size $L$ of the block marks a qualitative difference between 3D and 2D block tensor RG. Retaining a constant number of couplings in 3D loses its justification even off the criticality, which works in 2D due to the saturation of $S\to S_{0}$; this poses a threat to [Approximate $Z$ condition]. More details come later. Even if the block-tensor RG is exact and [Approximate $Z$ condition] is strictly satisfied, the scaling behavior of entanglement bodes ill for [RG flow condition]. Let’s say a block-tensor scheme is applied to the nearest-neighbor Ising model on a cubic lattice, how does the tensor RG flow look like? At infinite temperature, or $\beta=1/T=0$, after a gauge transformation [9], the tensor $A\to A^{\text{hi}}$ only has a single nonzero component, $A^{\text{hi}}_{000000}=1$; the entanglement entropy $S(\rho_{\vec{s}_{x+}})$ vanishes so $\alpha=F=0$ in Eq. (5). Now we consider $0<\beta\ll 1$. The entanglement $S(\rho_{\vec{s}_{x+}})$ of the resultant tensor has a small but non-vanishing $\alpha\neq 0$ in Eq. (5). A well-behaved RG transformation should generate an RG flow towards the high-temperature fixed point $A^{\text{hi}}$ with zero entanglement entropy, because the Ising model with $\beta\ll 1$ is in high-temperature phase. However, Eq. (5) shows that the coarser tensors generated by the block-tensor scheme has growing entanglement entropy due to the growth of $L$ and an $\alpha\neq 0$. Therefore, the 3D block-tensor RG, even if performed exactly by retaining infinite states, fails to satisfy [RG flow condition] in the vicinity of the high-temperature fixed point. Notice that the failure is not due to any approximation. Instead, as is indicated by the area law in Eq. (5), the incorrect RG flow is closely related to the violation of [UV-free condition]—the diverging term $\alpha L$ arises from microscopic details reflected in $\alpha$.
@@
The area law in Eq. (5) has an incarnation as a tensor-network toy model that mimics the entanglement structure of the high-temperature phase. The toy model captures the remnant entanglement as matrices sitting on the boundary. It is known in 2D as a corner-double line (CDL) structure [18]. We call its 3D generalization an edge-double-line (EDL) structure, since the boundaries between faces of a cube are edges, see Fig. 1; the EDL structure transforms under the exact block-tensor RG with rescaling factor $b$ as,
@@
Figure 2: Estimation of 3D Ising scaling dimensions $\Delta$ using high-order tensor renormalization group [29, 30] and its linearization [31]. Results are organized as per spin-flip $\mathbb{Z}_{2}$ symmetry. Dashed lines are best-known results [32]. Most approximation errors of $\sigma$ and $\epsilon$ operators lie between $10\%$ and $20\%$ and fail to decrease with more couplings retained.
@@
So far we have focused on the exact block-tensor RG in 3D and [Approximate $Z$ condition] is strictly satisfied. Next, we study the implication of the area law for a practical implementation of 3D block tensor, where we insist on keeping only $\chi$ states during RG. The failure of [UV-free condition] reflected in Eqs. (5) and (6) poses a challenge to 3D block-tensor RG as numerical techniques. Large amounts of computation power and memory are wasted on lattice-scale information $\alpha L$, while the finite and interesting universal information (for example, the 3D Ising universality class has $F=0.0623$ [33]) is buried among them. Therefore, for 3D block-tensor maps, increasing $\chi$ might not be a systematical way to approach the exact solution (see Fig. 2), in a striking contrast to the 2D beyond-simple-block TNRG results 111 See Fig. 3 in Ref. [19] and Table I in the supplementary material of Ref. [43]. . The numerical results in Fig. 2 can be reproduced using the python codes published in Ref. [35].
@@
Curiously, the failure of [RG flow condition] is hidden in numerical calculations due to a truncation effect. We can understand this truncation effect by feeding $A^{\text{EDL}}(e)$ into a 3D block-tensor RG with a fixed $\chi$. Assume without loss of generality that the edge-matrix is positive semi-definite and diagonal $e_{ij}=e_{i}\delta_{ij},e_{i}\geq 0$; the largest diagonal element is normalized to $e_{0}=1$, and the next largest one is $e_{1}<1$. We first perform the exact block-tensor RG with rescaling factor $b$, see Fig. 1 and Eq. (6). The truncation is based on the eigenvalue spectrum of a 1-face density matrix constructed from $A^{\prime}=A^{\text{EDL}}(e^{\otimes b})$ (recall that $A^{\prime}$ can be interpreted as a ground state wavefunction) according to Eq. (1). The largest eigenvalue is $e_{0}^{4b}=1$; the next largest one is $e_{1}$, with a degeneracy $4b$. As long as $\chi<4b$, following the often-adopted numerical practice of throwing away all states in the degenerate subspace, all states except the one with the largest eigenvalue are truncated; the edge-matrix $e$ is renormalized to $e^{\text{hi}}$. Therefore, under the approximated block-tensor RG, $A^{\text{EDL}}(e)\mapsto A^{\prime}=A^{\text{hi}}$ due to this truncation effect. The “seemingly correct RG flow” is observed in numerical practices [25, 26]. The caveat is that keeping a constant number of states leads to a bad approximation, which potentially undermines the validity of the 3D block-tensor RG. The seemingly satisfaction of [RG flow condition] is achieved at the cost of violating [Approximate $Z$ condition].
@@
Conclusion and remarks—The entanglement entropy area law is an impediment to 3D block transformation in tensor-network representation. A similar difficulty of a logarithmic growth of entanglement [5] arises in 2D criticality. Many advanced 2D TNRG schemes [18, 19, 20, 21, 22, 23, 24] applicable to criticality have been proposed, all striving to filter the remnant entanglement. Yet an entanglement-filtering-enhanced 3D TNRG scheme is still missing, even though the linear growth of entanglement in 3D systems, no matter critical or not, is worse. We must stress here that in 2D, the entanglement filtering was luxury for better accuracy, whereas it is basic necessity in 3D for any sensible RG. All existing 3D TNRG schemes [29, 36, 37, 38] are variations of the simple block-tensor idea. Entanglement filtering in 3D should be manipulation of the tensor network to renormalize out the pre-factor $\alpha$ of the 3D area law in Eq. (5); this indicates that the EDL structure in Fig. 1 provides a good guidance for the design of a better 3D TNRG. In Ref. [39], we propose a 3D entanglement-filtering scheme that can clear up the EDL structure, as well as a “sphere-like” structure [22]. Numerical results on the 3D Ising model suggest that the scheme in Ref. [39] manages to tame the area-law term in Eq. (5).
@@
Finally, we want to point out some implications for entropic c-theorems [40, 28]. In quantum field theories, the relationship between RG and entanglement entropy takes the form of various entropic c-theorems. The original c-theorem by Zamolodchikov [41] states that a central-charge-like $c$ quantity of 2D renormalizable quantum field theories is non-increasing along RG trajectories; $c$ is stationary at a fixed point and is proportional to the central charge of the corresponding conform field theory. Entropic c-theorems generalize the central-charge-like quantity to the universal contribution to the entanglement entropy of a sphere for a general $d$-dimensional quantum field theory, taking the value $F$ in Eq. (5) in 3D. In the context of an entropic c-theorem, the RG coarse graining is understood as increasing the radius of an entangling surface [28], in analogy to the block-tensor RG; the divergent terms in the entanglement entropy similarly haunts the analysis [42]. One may hope that the idea of entanglement filtering developed in numerical TNRG and its recent rigorous realization [8, 9] might inspire new ideas on how to deal with those divergent terms.
Acknowledgements.
@@
We thank Slava Rychkov for clarifying the conformal-field-theory prediction of 3D Ising critical exponents. We thank Raghav G. Jha and Zhiyuan Xie for explaining tensor-network RG analysis on 3D Ising and Potts model. We thank Satoshi Morita for helping with the parallel computing implementation of 3D tensor-network RG. We thank Shumpei Iino, Wei-Lin Tu, Katsuya Akamatsu, Kenji Homma, Antoine Tilloy, Takeo Kato, Hirokazu Tsunetsugu, Yuan Yao, Weiguang Cao and Shihwen Hor for fruitful discussions. X.L. is grateful to the support of the Global Science Graduate Course (GSGC) program of the University of Tokyo. This work is financially supported by MEXT Grant-in-Aid for Scientific Research (B) (23H01092). The computation in this work has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.

### v3

@@ 5.2 Exact 3D block-tensor map
The advantage of TNRG in 2D makes it a promising candidate as a 3D real-space RG. One thing to note is that, without the viewpoint of entanglement entropy and its area laws, whether the block-tensor RG in 3D is better or worse than that in 2D is not straightforward. Recall that in Wilson’s momentum-shell RG, systems in dimensions $d>4$ are easy, where mean-field treatment of Landau-Ginzburg field is enough; but in $d<4$, perturbative $\epsilon$-expansion is necessary and $d=2$ is harder than $d=3$ since $\epsilon=4-d$ is larger. Interestingly, numerical experiments [wang2014, jha-3dpotts] would seemingly imply that block-tensor RG in 3D improves [RG flow condition] compared to 2D. However, the truncation effects, as a cause of those observations, become more transparent through the lens of entanglement entropy.
@@
Figure 2: (Upper left) A tensor $A^{\text{EDL}}(e)$ with an EDL structure is the tensor product of 12 copies of edge-matrix $e$ located at 12 edges of a cube block; every 4 matrix-indices on the same face are grouped together to be a thicker leg of the 6-leg $A^{\text{EDL}}(e)$ . (Lower right) A block-tensor RG with rescaling factor $b=2$. A total of $12\times 8=96$ copies of edge-matrix $e$ form 3 types: (i) Around each of the 6 axes inside the block, 4 copies of $e$ form a loop and becomes a number, so there are $4\times 6$ copies; (ii) On each of 6 faces, copies of $e$ form 4 entanglement pairs inside the face, so there are $2\times 4\times 6$ copies; (iii) The remaining $2\times 12$ copies sit on the 12 edges of the larger block, each edge having 2 copies side by side. Only type-(iii) copies survive the exact block-tensor RG; the other two types become an overall multiplication factor. The coarser tensor has an EDL structure built from $e\otimes e$: $A^{\prime}=A^{\text{EDL}}(e\otimes e)$.
@@
The assessment of the block-tensor RG in 3D parallels its analysis in 2D. We have a 6-leg tensor $A_{\vec{s}_{x+},\vec{s}_{x-},\vec{s}_{y+},\vec{s}_{y-},\vec{s}_{z+},\vec{s}_{z-}}$ (for example, $x+$ denotes the leg along the positive $x$ axis), corresponding to a cubic block of spins. The TNRG map in Eq. (5) satisfies $Z_{N}(A)=Z_{N/b^{3}}(A^{\prime})$. For a large block with linear size $L$, the area law of 3D quantum field theories says that the entanglement entropy of a subsystem on a face of the cube, $S(\rho_{\vec{s}_{x+}})$ for example, scales like [kitaev-preskill, nishioka2018]
@@
where $\alpha$ contains microscopic details and $F$ encodes universal properties of the system. This entropy should not be confused with the mutual information between the inside and the outside of the cube mediated through the face. Rather, it is from the correlations between the two surface areas mediated through the edge region. The area law in Eq. (6) reveals a lot about 3D block-tensor transformation. The growth of $S(\rho_{\vec{s}_{x+}})$ with the size $L$ of the block marks a qualitative difference between 3D and 2D block tensor RG. Retaining a constant number of couplings in 3D loses its justification even off the criticality, which works in 2D due to the saturation of $S\to S_{0}$; this poses a threat to [Approximate $Z$ condition]. More details come later. Even if the block-tensor RG is exact and [Approximate $Z$ condition] is strictly satisfied, the scaling behavior of entanglement bodes ill for [RG flow condition]. Let’s say a block-tensor scheme is applied to the nearest-neighbor Ising model on a cubic lattice, how does the tensor RG flow look like? At infinite temperature, or $\beta=1/T=0$, after a gauge transformation [kennedy-lowT], the tensor $A\to A^{\text{hi}}$ only has a single nonzero component, $A^{\text{hi}}_{000000}=1$ [^2] For this high-$T$ fixed-point tensor, the index value 0 means an equal-weight superposition of all the spin-based states represented by this leg. For example, if a leg of the initial tensor corresponds to a single spin, then the state 0 represents $(|\uparrow\rangle+|\downarrow\rangle)/\sqrt{2}$. This also applies to the high-$T$ edge matrix $e^{\text{hi}}$ in the discussion below Eq. (7). ; the entanglement entropy $S(\rho_{\vec{s}_{x+}})$ vanishes so $\alpha=F=0$ in Eq. (6). Now we consider $0<\beta\ll 1$. The entanglement $S(\rho_{\vec{s}_{x+}})$ of the resultant tensor has a small but non-vanishing $\alpha\neq 0$ in Eq. (6). A well-behaved RG transformation should generate an RG flow towards the high-temperature fixed point $A^{\text{hi}}$ with zero entanglement entropy, because the Ising model with $\beta\ll 1$ is in high-temperature phase. However, Eq. (6) shows that the coarser tensors generated by the block-tensor scheme has growing entanglement entropy due to the growth of $L$ and an $\alpha\neq 0$. Therefore, the 3D block-tensor RG, even if performed exactly by retaining infinite states, fails to satisfy [RG flow condition] in the vicinity of the high-temperature fixed point. Notice that the failure is not due to any approximation. Instead, as is indicated by the area law in Eq. (6), the incorrect RG flow is closely related to the violation of [UV-free condition]—the diverging term $\alpha L$ arises from microscopic details reflected in $\alpha$. At criticality, the growth of the entanglement entropy suggests that an exact 3D block-tensor RG cannot exhibit a critical fixed-point tensor, since the entanglement entropy is a single-number characterization of the tensor. This bodes ill for the hope of obtaining a critical fixed-point tensor numerically using a 3D block-tensor map [^3] See Ref. [Lyu:Kawashima:2024] for numerical evidence. .
@@
The area law in Eq. (6) has an incarnation as a tensor-network toy model that mimics the entanglement structure of the high-temperature phase. The toy model captures the remnant entanglement as matrices sitting on the boundary. It is known in 2D as a corner-double line (CDL) structure [gu-wen2009]. We call its 3D generalization an edge-double-line (EDL) structure, since the boundaries between faces of a cube are edges, see Fig. 2; the EDL structure transforms under the exact block-tensor RG with rescaling factor $b$ as,
@@
5.3 Numerical 3D block-tensor map
@@
So far we have focused on the exact block-tensor RG in 3D and [Approximate $Z$ condition] is strictly satisfied. Next, we study the implication of the area law for a practical implementation of 3D block tensor, where we insist on keeping only $\chi$ states during RG. The failure of [UV-free condition] reflected in Eqs. (6) and (7) poses a challenge to 3D block-tensor RG as numerical techniques. Large amounts of computation power and memory are wasted on lattice-scale information $\alpha L$, while the finite and interesting universal information (for example, the 3D Ising universality class has $F=0.0623$ [f3dising]) is buried among them.
The linear growth of the entanglement entropy in 3D indicates a rapid growth of RG truncation errors using a block-tensor map, since the size of the block grows exponentially with the RG step $n$. This expectation can be checked by plotting RG truncation errors with respect to the RG step. We apply the higher-order tensor renormalization group (HOTRG) [hotrg2012] to the 3D Ising model with the nearest-neighbor interaction at the estimated critical temperature. A single RG step in the HOTRG consists of three collapses in $x,y$ and $z$ directions, each collapse having two truncation errors. In total, there are $2\times 3=6$ truncation errors for a single RG step. We use the maximum among these six errors to represent the truncation error of a single RG step, and plot how it changes with the RG step $n$ at bond dimensions $\chi\leq 10$ in Fig. 3. The block-tensor map in 3D loses its justification as a reliable real-space RG transformation due to this rapid growth of the RG truncation errors [^4] The HOTRG [hotrg2012] may sometimes seem to produce estimates of scaling dimensions with relatively high accuracy for a “magic” bond dimension $\chi=14$. We want to point out that the seemingly high accuracy could be an accident that is not robust against the change of the bond dimension. This accidental high accuracy is also observed in our calculations in Fig. 5, where the relative error of the energy-density scaling dimension $\Delta_{\epsilon}$ is smaller than $10^{-2}$ at $\chi=12$. .
Figure 3: The rapid growth of the RG error with respect to the RG step. The temperature is set to be the estimated critical temperature obtained by studying the RG flows, as is explained in Refs. [lyu-kawashima2021, Lyu:Kawashima:2024]. The RG error grows to more than $10\%$ just after the first RG step. Near the critical fixed point, which is around RG steps $3\leq n\leq 5$, the RG error increases from $10\%$ to $30\%$ when the bond dimension increases from 44 to $10$. At the largest bond dimension we reach, $\chi=22$ (not shown in the Figure), the RG truncation error near the critical fixed point has grown to about $38\%$.
Furthermore, as a single-number characterization of the tensor, the linear growth of the entanglement entropy suggests that it would be unreasonable to expect the block-tensor RG transformation to exhibit a critical fixed-point tensor. This poses difficulty in carrying out the Wilsonian RG prescription in a 3D block-tensor RG, where a fixed point is first identified and then the scaling dimensions can be estimated from the linearized RG map at the fixed point (see a short review of the linearized RG method for the HOTRG in Appendix A). The growth of the RG error shown in Fig. 3 demonstrates that the tensor is not fixed near the supposed critical “fixed point” [^5] More numerical evidence of the fact that the “fixed-point” tensor is not fixed can be found in Ref. [Lyu:Kawashima:2024]. . The consequence is that the estimates of the scaling dimensions using the linearized RG map fail to converge with respect to the RG step (see Fig. 4). Apart from the large RG error near the critical “fixed point” shown in Fig. 3, the drifting of the estimates of the scaling dimensions in Fig. 4 is another reason that a 3D block-tensor map (like the HOTRG) becomes an unreliable real-space RG method.
Figure 4: Drifting of the estimates of scaling dimensions that happens at most bond dimensions $\chi\leq 22$. The example shown here is at $\chi=8$. The spin and energy density operators are denoted as $\sigma,\epsilon$, respectively, while $\partial_{i}\sigma,\partial_{i}\partial_{j}\sigma$ and $\partial_{i}\epsilon$ are their descendant operators. The energy-momentum operator is denoted as $T_{ij}$. The estimate of $\epsilon$ drifts with respect to the RG step.
The increase of the RG error with the bond dimension in Fig. 3 suggests that increasing $\chi$ might not be a systematic way to approach the exact solution in practical numerical calculations. However, the drifting of the estimates of scaling dimensions in Fig. 4 makes the study of the dependency of the estimates on $\chi$ trickier. One way to decide which RG step to choose for linearizing the RG map is by examining the difference of adjacent tensors and choosing the step in which the tensor is the most stable. Using this strategy [^6] We also tried fixing the RG step for studying the $\chi$ dependency; just like Fig. 5, we failed to see any systematic trend of the estimates of the scaling dimensions when $\chi$ increases. , we plot the $\chi$ dependency of the scaling dimensions in Fig. 5. This failure of a systematic improvement in a 3D block-tensor map when the bond dimension $\chi$ increases is in a striking contrast to the 2D beyond-simple-block TNRG results [^7] See Fig. 3 in Ref. [tnr2015] and Table I in the supplementary material of Ref. [evenbly-vidal-local]. . The numerical results in Figs. 3 to 5 can be reproduced using the python codes published in Ref. [Lyu:bkten3d:codes].
Figure 5: Estimation of 3D Ising scaling dimensions $\Delta$ using high-order tensor renormalization group [hotrg2012, hotrg-paral2022] and its linearization [lyu-kawashima2021]. Results are organized as per spin-flip $\mathbb{Z}_{2}$ symmetry. Dashed lines are results from conformal bootstrap [bootstrap-3dising]. Most approximation errors of $\sigma$ and $\epsilon$ operators lie between $10\%$ and $20\%$ and fail to decrease with more couplings retained.
@@
Curiously, the failure of [RG flow condition] is hidden in numerical calculations due to a truncation effect. We can understand this truncation effect by feeding $A^{\text{EDL}}(e)$ into a 3D block-tensor RG with a fixed $\chi$. Assume without loss of generality that the edge-matrix is positive semi-definite and diagonal $e_{ij}=e_{i}\delta_{ij},e_{i}\geq 0$; the largest diagonal element is normalized to $e_{0}=1$, and the next largest one is $e_{1}<1$. We first perform the exact block-tensor RG with rescaling factor $b$, see Fig. 2 and Eq. (7). The truncation is based on the eigenvalue spectrum of a 1-face density matrix constructed from $A^{\prime}=A^{\text{EDL}}(e^{\otimes b})$ (recall that $A^{\prime}$ can be interpreted as a ground state wavefunction) according to Eq. (1). The largest eigenvalue is $e_{0}^{4b}=1$; the next largest one is $e_{1}$, with a degeneracy $4b$. As long as $\chi<4b$, following the often-adopted numerical practice of throwing away all states in the degenerate subspace, all states except the one with the largest eigenvalue are truncated; the edge-matrix $e$ is renormalized to $e^{\text{hi}}$. Therefore, under the approximated block-tensor RG, $A^{\text{EDL}}(e)\mapsto A^{\prime}=A^{\text{hi}}$ due to this truncation effect. The “seemingly correct RG flow” is observed in previous numerical practices [wang2014, jha-3dpotts], as well as in the numerical calculation performed in this paper. The caveat is that keeping a constant number of states leads to a bad approximation, which potentially undermines the validity of the 3D block-tensor RG. The seemingly satisfaction of [RG flow condition] is achieved at the cost of violating [Approximate $Z$ condition].
6 Conclusion and remarks
We have demonstrated that the entanglement-entropy area laws offer a unified framework for understanding the properties of the block-tensor map in TNRG. They indicate how large the RG approximation errors are and how they change with respect to the RG step. They help diagnose whether the block-tensor map leads to correct RG flows, or specifically, whether it exhibits any critical fixed-point tensor. Moreover, it indicates the location of microscopic information, and thus can help guide the design of a more efficient TNRG map.
@@ 6 Conclusion and remarks
In 3D, the entanglement-entropy area law is an impediment to a block transformation in tensor-network representation. A similar difficulty of a logarithmic growth of entanglement [vidal2003] arises in 2D criticality. Many advanced 2D TNRG schemes [gu-wen2009, tnr2015, tnr-plus2017, loop-tnr2017, gilt2018, harada2018, fet2018] applicable to criticality have been proposed, all striving to filter the remnant entanglement. Yet an entanglement-filtering-enhanced 3D TNRG scheme is still missing, even though the linear growth of entanglement in 3D systems, no matter critical or not, is worse. After the proposal of the first 3D block-tensor scheme, the higher-order tensor renormalization group (HOTRG) [hotrg2012], many attempts to develop the 3D TNRG [atrg2020, triadtrg2019, Kadoh_2022] focus on reducing its computational complexity. However, it is not a higher computational complexity [^8] The HOTRG has computational complexity $O(\chi^{7})$ in 2D and $O(\chi^{11})$ in 3D. but the overflow of the microscopic entanglement that makes the 3D problem qualitatively different from the 2D one. We must stress here that in 2D, the entanglement filtering was luxury for better accuracy, whereas it is basic necessity in 3D for any apt TNRG scheme. Entanglement filtering in 3D should be manipulation of the tensor network to renormalize out the pre-factor $\alpha$ of the 3D area law in Eq. (6); this indicates that the EDL structure in Fig. 2 provides a good guidance for the design of a better 3D TNRG. In Ref. [Lyu:Kawashima:2024], we propose a 3D entanglement-filtering scheme that can clear up the EDL structure, as well as a “sphere-like” structure [gilt2018]. Numerical results on the 3D Ising model suggest that the scheme in Ref. [Lyu:Kawashima:2024] manages to tame the area-law term in Eq. (6).
@@
Finally, we want to point out some implications for entropic c-theorems [casini2007, nishioka2018]. In quantum field theories, the relationship between RG and entanglement entropy takes the form of various entropic c-theorems. The original c-theorem by Zamolodchikov [zomo-ctheorem] states that a central-charge-like $c$ quantity of 2D renormalizable quantum field theories is non-increasing along RG trajectories; $c$ is stationary at a fixed point and is proportional to the central charge of the corresponding conform field theory. Entropic c-theorems generalize the central-charge-like quantity to the universal contribution to the entanglement entropy of a sphere for a general $d$-dimensional quantum field theory, taking the value $F$ in Eq. (6) in 3D. In the context of an entropic c-theorem, the RG coarse graining is understood as increasing the radius of an entangling surface [nishioka2018], in analogy to the block-tensor RG; the divergent terms in the entanglement entropy similarly haunts the analysis [casini2015]. One may hope that the idea of entanglement filtering developed in numerical TNRG and its recent rigorous realization [kennedy-hiT, kennedy-lowT] might inspire new ideas on how to deal with those divergent terms.
[…]
@@
Acknowledgements We thank Slava Rychkov for clarifying the conformal-field-theory prediction of 3D Ising critical exponents. We thank Raghav G. Jha and Zhiyuan Xie for explaining tensor-network RG analysis on 3D Ising and Potts model. We thank Satoshi Morita for helping with the parallel computing implementation of 3D tensor-network RG. We thank Shumpei Iino, Wei-Lin Tu, Katsuya Akamatsu, Kenji Homma, Antoine Tilloy, Takeo Kato, Hirokazu Tsunetsugu, Yuan Yao, Weiguang Cao and Shihwen Hor for fruitful discussions. X.L. is grateful to the support of the Global Science Graduate Course (GSGC) program of the University of Tokyo. The peer reviewing process of this work was done after X.L. moved to IHES. This work is financially supported by JSPS KAKENHI Grant Number 23K25789. The computation in this work has been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.
Declarations
[…]
Competing interests The authors have no competing interests to declare that are relevant to the content of this article.
[…]
Code availability The numerical results in this work can be reproduced using the codes published in the Ref. [Lyu:bkten3d:codes].
Appendix A Linearized RG map of the HOTRG
In the context of TNRG, the Wilsonian RG prescription for extracting scaling dimensions from the linearization of the RG map has been developed in Refs. [lyu-kawashima2021, Ebel:2024LDO]. In this appendix, we demonstrate how to perform this prescription for the 2D HOTRG. The generalization to the 3D HOTRG is straightforward and has been expounded in Ref. [lyu:kawashima:2025lattsym]. The RG transformation of the 2D HOTRG [hotrg2012] is a nonlinear map $A\mapsto A^{\prime}=\mathcal{T}(A)$, which is a composition of the following two collapses in two directions,
[…]
The order of the collapse is chosen arbitrarily to be $y\rightarrow x$ here. The triangular tensors $p_{x},p_{y}$ are known as isometric tensors and can be determined using either the higher-order singular value decomposition [hotrg2012], or, equivalently, the projective truncations [Evenbly:tnralgo, lyu:kawashima:2025lattsym].
In the first stage of the Wilsonian RG prescription, by tuning the temperature of the model, one determines a fixed-point tensor $A_{*}$ of the above RG map such that it is invariant under the RG map: $A_{*}=\mathcal{T}(A_{*})$. Due to the gauge freedom of a tensor-network representation of the partition function, a gauge-fixing scheme is necessary at this stage [lyu-kawashima2021, Ebel:2025Rot] for obtaining a fixed-point tensor $A_{*}$ in numerical calculations. The extent to which the tensor $A_{*}$ is fixed numerically can be greatly improved by the Newton Method [Ebel:2025Rot].
In the second stage, when $A_{*}$ has been determined, let us denote the corresponding isometric tensors to be $p_{x*},p_{y*}$. The linearized map of the HOTRG is a linear map $\delta A\mapsto\delta A^{\prime}=\mathcal{R}(\delta A)$, which can be built from $A_{*},p_{x*},p_{y*}$ according to Refs. [lyu-kawashima2021, Ebel:2024LDO] as the following composition:
[…]
where
[…]
According to Wilsonian RG theory, the scaling dimensions $\{\Delta_{i}\}$ of the criticality can be extracted from the eigenvalues $\{\lambda_{i}\}$ of the linearized RG map $\mathcal{R}$ in Eq. (9):
$b^{d-\Delta_{i}}=\lambda_{i},$ (10)
where $b=2$ is the rescaling factor of the RG transformation and $d=2$ is the spatial dimensionality of the system.